OpenStudy (anonymous):

A 12-foot ladder is leaning against a house with the base of the ladder 5 feet from the house. How high up the house does the ladder reach? If necessary, round to the nearest tenth foot.

OpenStudy (anonymous):

Use pythagorean theorem

OpenStudy (anonymous):



OpenStudy (anonymous):

i need help on that

OpenStudy (anonymous):

In a RIGHT TRIANGLE, \[a^2+b^2=c^2\] where c is the hypotenuse, opposite the right angle.

OpenStudy (anonymous):

\[12^2-5^2=x^2\]

OpenStudy (anonymous):

\[x=\sqrt{12^2-5^2}=\sqrt{119} \approx 10.9ft\]